OpenStudy (anonymous):

Find the derivative of f(x) = -7/x at x = -3

OpenStudy (sidsiddhartha):

f(x)=-7/x so write x in the form of (x)^-1 and 7 is a constant so d/dx(fx)=-7*(d/dx)(x^-1)=-7*(-1*x^-2)=7/x^2 at x-=3 derivative=7/9

OpenStudy (anonymous):

thanks you!

OpenStudy (sidsiddhartha):

wlcm :)